# 3DGS-CD: 3D Gaussian Splatting-based Change Detection for Physical Object Rearrangement

### v1

@@ Abstract
We present 3DGS-CD, the first 3D Gaussian Splatting (3DGS)-based method for detecting physical object rearrangements in 3D scenes. Our approach estimates 3D object-level changes by comparing two sets of unaligned images taken at different times. Leveraging 3DGS’s novel view rendering and EfficientSAM’s zero-shot segmentation capabilities, we detect 2D object-level changes, which are then associated and fused across views to estimate 3D changes. Our method can detect changes in cluttered environments using sparse post-change images within as little as 18s, using as few as a single new image. It does not rely on depth input, user instructions, object classes, or object models – An object is recognized simply if it has been re-arranged. Our approach is evaluated on both public and self-collected real-world datasets, achieving up to 14% higher accuracy and three orders of magnitude faster performance compared to the state-of-the-art radiance-field-based change detection method. This significant performance boost enables a broad range of downstream applications, where we highlight three key use cases: object reconstruction, robot workspace reset, and 3DGS model update. Our code and data will be made available at https://github.com/520xyxyzq/3DGS-CD.
@@ I INTRODUCTION
3D change detection involves identifying changed objects or regions in the environment from two sets of local observations taken at different times or from new observations of a previously modeled scene. It is a critical task in robotics as it can accommodate not only short-term, fully observed scene dynamics but also long-term scene changes where the transition process is typically unobserved. Such long-term changes are particularly important for robot operations, as environmental changes often occur without being fully noticed or monitored.
@@
In our work, we leverage 3DGS as scene representation to identify 3D scene changes including object removal, insertion and movement from multi-view images. We exploit the zero-shot segmentation capability of EfficientSAM [11] to compare pre- and post-change images at the same viewpoints, associating and fusing the 2D object changes to obtain accurate 3D object segments and pose changes.
Our method has the following key advantages: (1) It can handle sparse post-change image inputs, requiring as few as a single new image to detect 3D changes. (2) It requires no depth sensors or monocular depth estimators. (3) It does not rely on pre-defined object classes, models or object detectors – An object is recognized simply if it has been moved, removed or inserted. (4) It requires no user instructions such as user-provided click or language prompts.
@@ II-A Change detection from RGB images
2D change detection from image pairs has been a long-studied problem. Most existing methods assume minimal viewpoint differences between images (e.g. [12]). While some approaches can handle larger viewpoint shifts (e.g. [13]), they are often restricted to the provided views and cannot fully capture physical changes in 3D.
Traditionally, 3D change detection from multi-view RGB images has relied on hand-crafted techniques to identify changes from unaligned images [4, 5, 6]. NeRF-based methods have been explored to directly compare pre- and post-change images from the same viewpoints, but only to a limited extent. NeRF-Update [9] combines NeRF with the Segment Anything Model (SAM) [14] to detect object movements, while C-NeRF [10] models scene changes as direction-consistent radiance differences between two NeRFs. However, these methods are limited by NeRF’s slow rendering speed. In contrast, we leverage 3DGS’s real-time rendering capability for efficient 3D change detection.
@@
3DGS has recently emerged as a key scene representation in robotics but it lacks native support for representing scene changes. While dynamic 3DGS methods (e.g. [15]) have been developed to capture short-term dynamics in video sequences, these methods typically require continuous observation of the dynamic process. Our method addresses this gap by detecting changes in 3DGS-represented scenes, offering a more robust solution for capturing 3D scene dynamics.
@@ III-B EfficientSAM
The segment anything model (SAM) [14] is a foundation model for image segmentation from input prompts such as boxes or points. EfficientSAM[11] is a light-weight SAM model enabled by masked image pretraining. EfficientSAM adopts an encoder-decoder architecture, i.e. $\mathcal{S}=\mathcal{S}_{e}\circ\mathcal{S}_{d}$. The encoder $\mathcal{S}_{e}$ predicts the embedding $f\in\mathbb{R}^{h\times w\times d}$ for input image $I\in\mathbb{R}^{H\times W\times 3}$, and the decoder $S_{d}$ takes the image embedding $f$ and the prompt $\mathcal{P}$ to output the segmentation mask and the prediction confidence.
@@ IV-A Pre-change 3DGS training
We first train a 3DGS model $\mathcal{G}$ on the pre-change images $\{I_{m}\}_{m=1}^{M}$. Following the standard 3DGS training procedure [8], we use a structure-from-motion (SfM) algorithm [16] to estimate pre-change camera parameters and generate a sparse point cloud, initializing 3D Gaussians for optimization. 111 For the success of SfM, we typically need the pre-change images to sufficiently cover the pre-change scene. But this requirement can be relaxed with sparse-view 3DGS training techniques, e.g. InstantSplat [17].
@@ IV-C EfficientSAM embedding for 2D change detection
(a) Rendered
(b) Captured
(c) Chg. msk
(d) move-out
(e) move-in
Figure 2: 2D object-level change detection at post-change views: Rendered and captured images $\rightarrow$ Embedding $\rightarrow$ Cosine similarity $\rightarrow$ Change map $\rightarrow$ Object masks. Despite incomplete 2D object masks due to occlusions or prediction failures, we can fuse the multi-view masks to obtain more complete object templates and classify object changes (Sec. IV-D).
We first coarsely detect 2D changes with EfficientSAM [11] image embeddings (Fig. 2(a-c)). The SAM [14] image embedding, which encodes rich semantic information, has been shown to enable robust zero-shot change detection [9, 18]. However, probing SAM’s latent space can be computationally expensive. We therefore adopt the lightweight EfficientSAM encoder to extract image embeddings on the rendered and captured post-change images, and threshold their cosine similarity maps to obtain the change masks:
@@
As shown in Fig. 2 (d-e), the high-confidence 2D segments on the rendered images are used as move-out masks (previous location) for the re-arranged objects, whereas the high-confidence masks on the captured images are identified as object move-in masks (new location):
@@ IV-D Change-aware object association
Starting from the object move-out masks on the first view, we iteratively associate the segments across remaining views by matching the partial point clouds. Specifically, we use the Hungarian algorithm to perform nearest-neighbor matching between the current object point cloud and the next segment with the Hausdorff distance. The matched new segments are added to the object templates while the unmatched segments are used to initialize new object templates. We also verify the semantic similarity of the matched segments by comparing their EfficientSAM embeddings. If their cosine similarity is below 40%, a new object template is created. After the association is done, we obtain the object templates as the union of multi-view point clouds and features:
@@ IV-F Object pose change estimation
For moved objects, we use the 2D-3D correspondences from Sec.IV-D, and apply the RANSAC-PnP algorithm to coarsely estimate the object pose changes. The initial estimates are further optimized in Sec. IV-G.
@@
Figure 3: Estimated 3D object masks projected on novel views. We show the projected object move-out (previous location) and move-in (new location) masks on the pre-change 3DGS render (left) and the post-change capture (right) from the same novel viewpoint for each scene. Please check out the supplementary video for mask projections on more views.
(a) Ground truth
(b) Ours
(c) MTP[21]
@@ V-C Baselines
D-NeRF-Diff [24]: Another baseline by C-NeRF authors that trains a D-NeRF with the two image sets as two time steps. It computes change masks using the aboslute color differences between evaluation-view D-NeRF renders.
On the 3DGS-CD dataset, we use MTP as baseline to predict change masks on rendered and captured images at evaluation views  [^3] C-NeRF is not used as a baseline on 3DGS-CD since it fails to train a post-change NeRF from the sparse post-change images.. We adopt the ViT-B+RVSA model fine-tuned on the CDD dataset [25] as it performs best on our data. On the NeRF-CD dataset, we compare our method against C-NeRF, C-NeRF-Diff and D-NeRF-Diff.
@@ V-D Results
Our method is more accurate than the MTP baseline by a large margin, although it underperforms in runtime due to the inference efficiency of neural networks. We acknowledge that MTP’s prediction accuracy could be improved through fine-tuning on similar data. However, our method exhibits strong zero-shot performance without the need for in-domain fine-tuning.
@@
TABLE II: Quantitative evaluation of our method on the test scene Cats from the NeRF-CD [10] dataset.
@@ VI-A Object removal as prompt for object reconstruction
To illustrate this, we use the SOTA prompt-based 3DGS segmentation method SAGA [26] to isolate an target object from its 3DGS-represented scene, as shown in Fig. 6. While SAGA explicitly accounts for multi-level granularity, it still struggled to segment the entire object with a single click prompt ( or ), even with the best thresholds. We had to provide multiple precise clicks ( ) to extract the complete object. However, providing such precise clicks can be difficult on mobile devices. Similarly, using a language prompt generated by GPT-4o also failed to isolate the whole object.
@@
To address this, we propose object removal as an unambiguous prompt, as shown in Fig. 7. After capturing images of the object in its scene, users only have to remove the object, and take a few more images for the object-removed scene. Our method can then consistently identify the target object from the pre-change images, allowing us to train an object-3DGS without ambiguity. Compared to the click or language prompts, the object removal prompt is more intuitive and requires minimal user efforts.
@@ VI-B Robot workspace reset
Figure 8: 3DGS-based change detection enables physical world reset with a robotic arm. Please check out the supp. material for the robot reset video.
@@
The physical world reset process can be summarized as follows: (1) RGB video scan to capture the initial state of the robot workspace; (2) Move objects in the workspace; (3) Capture a few more RGB images for the changed state of the workspace; (4) Use our method to estimate the moved objects’ shapes and pose changes in under 30s; (5) Robot pick and place to reset the state of the scene.
Key advantages of our physical world reset method over prior approaches (e.g. SayCan [27]) include: (1) No depth sensor or mono depth estimator is required; (2) No need for object class, model or object detectors: (3) No need for potentially ambiguous language prompts: (4) Precise reset of the object position and orientation.
Figure  8 shows a simulated experiment we conducted in PyBullet [28], where a robot arm resets the pose of a mug on a table. In this experiment, we used forward kinematics, rather than structure-from-motion (SfM) or visual localization, to estimate the pre- and post-change poses of the robot’s RGB camera. Without the SfM point cloud, the pre-change 3D Gaussians are randomly initialized before optimization. Once the middle mug’s pose change and shape are computed, we determine the pick-and-place poses for the robot gripper and use inverse kinematics to compute the joint motions. For simplicity, we use the initial gripper orientation as the grasp orientation and use the object’s centroid as the grasp position.
@@ VI-C Fast sparse-view-guided 3DGS update
In our method, we transform the in-object pre-optimized Gaussians according to the estimated object movements, and freeze all pre-trained Gaussian parameters. We then duplicate the Gaussians near the newly un-occluded regions and optimize only the duplicates to fill in the missing parts. Since these Gaussians are usually similar in position and color to the missing ones, they provide a good initialization for subsequent 3DGS optimization under sparse-view supervision. The Gaussians are optimized for 1000 step, during which we use the adaptive control strategy [8] to split, duplicate, reset and cull Gaussians every 100 steps. A visualization of the 3DGS optimization process is available in our supplementary video.
We test our method on the Soda scene from the NeRF-Update dataset [9], as shown in Fig. 9. In this scene, 174 pre-change images are used to train a 3DGS and 4 post-change images serve as guidance for the update. We evaluate our method on reconstruction quality for the soda can’s move-in and move-out regions with PSNR and SSIM, and runtime performance, as reported in Tab. III. Our method achieves comparable reconstruction quality and superior runtime performance compared to the NeRF-update [9] method. We also compare it with the SOTA sparse-view 3DGS training method, InstantSplat [8], which is only trained on the 4 post-change images. While InstantSplat is faster, it delivers lower reconstruction quality.

### v2

@@ Abstract
We present 3DGS-CD, the first 3D Gaussian Splatting (3DGS)-based method for detecting physical object rearrangements in 3D scenes. Our approach estimates 3D object-level changes by comparing two sets of unaligned images taken at different times. Leveraging 3DGS’s novel view rendering and EfficientSAM’s zero-shot segmentation capabilities, we detect 2D object-level changes, which are then associated and fused across views to estimate 3D change masks and object transformations. Our method can accurately identify changes in cluttered environments using sparse (as few as one) post-change images within as little as 18s. It does not rely on depth input, user instructions, pre-defined object classes, or object models – An object is recognized simply if it has been re-arranged. Our approach is evaluated on both public and self-collected real-world datasets, achieving up to 14% higher accuracy and three orders of magnitude faster performance compared to the state-of-the-art radiance-field-based change detection method. This significant performance boost enables a broad range of downstream applications, where we highlight three key use cases: object reconstruction, robot workspace reset, and 3DGS model update. Our code and data will be made available at https://github.com/520xyxyzq/3DGS-CD.
@@ I INTRODUCTION
3D change detection involves identifying changed objects or regions in the environment from two sets of local observations taken at different times or from new observations of a previously modeled scene. It is a critical task in robotics as it can accommodate not only short-term, continuously observed scene dynamics but also long-term scene changes where the transition process is typically unobserved. Capturing such long-term changes is particularly important for robot operations, as environmental changes often occur without being noticed or monitored.
@@
In our work, we leverage 3DGS as scene representation to identify 3D scene changes including object removal, insertion and movement from multi-view images. We exploit the zero-shot segmentation capability of EfficientSAM [11] to compare pre- and post-change images at the same viewpoints, associating and fusing the 2D object changes to obtain accurate object 3D change masks and 6D pose changes.
Our method has the following key advantages: (1) It can handle sparse post-change image inputs, requiring as few as a single new image to detect 3D changes. (2) It requires no depth sensors or monocular depth estimators. (3) It does not rely on pre-defined object classes, models or object detectors – An object is recognized simply if it has been moved, removed or inserted. (4) It requires no user instructions such as user-provided click or language prompts. (5) It is agnostic to the choice of radiance field models and could, in principle, work with any, such as NeRF[7] and 2DGS [12].
@@ II-A Change detection from RGB images
2D change detection from image pairs has been a long-studied problem. Most existing methods assume minimal viewpoint differences between images [13, 14, 15, 16, 17, 18, 19]. While some approaches can handle larger viewpoint shifts [20, 21, 22], they are often restricted to the provided views and cannot fully capture physical changes in 3D.
Traditionally, 3D change detection from multi-view RGB images has relied on hand-crafted techniques to identify changes from unaligned images [4, 5, 6]. NeRF-based methods have been explored to directly compare pre- and post-change images from the same viewpoints, but only to a limited extent. NeRF-Update [9] combines NeRF with the Segment Anything Model (SAM) [23] to detect object movements, while C-NeRF [10] models scene changes as direction-consistent radiance differences between two NeRFs. However, these methods are limited by NeRF’s slow rendering speed. In contrast, 3DGS offers real-time rendering speeds through its rasterization-based rendering pipeline. While early works such as SplatPose & Detect [24] have explored the use of 3DGS for anomaly detection, there has been little effort to directly apply 3DGS to the task of change detection. In this work, we leverage 3DGS’s efficient novel-view rendering capability for fast 3D scene change detection.
@@
3DGS has recently emerged as a key scene representation in robotics but it lacks native support for change representation. While dynamic 3DGS methods [25, 26, 27, 28, 29, 30, 31, 32] can capture short-term dynamics in video sequences, many of them typically rely on continuous spatio-temporal observation of the transformation process. Our method addresses this gap by detecting changes in 3DGS-represented scenes, offering a more robust solution for capturing 3D scene dynamics.
@@ III-B EfficientSAM
The segment anything model (SAM) [23] is a foundation model for image segmentation from input prompts such as boxes or points. EfficientSAM[11] is a light-weight SAM model enabled by SAM-leveraged masked image pretraining. EfficientSAM adopts an encoder-decoder architecture, i.e. $\mathcal{S}=\mathcal{S}_{e}\circ\mathcal{S}_{d}$. The encoder $\mathcal{S}_{e}$ predicts the embedding $f\in\mathbb{R}^{h\times w\times d}$ for input image $I\in\mathbb{R}^{H\times W\times 3}$, and the decoder $S_{d}$ takes the image embedding $f$ and the prompt $\mathcal{P}$ to output the segmentation mask and the prediction confidence.
@@
Figure 2: EfficientSAM embedding for 2D change detection at a post-change view (Sec. IV-C). While the 2D object masks may be incomplete or missing due to occlusions or prediction failures, associating and fusing masks from multiple views (Sec. IV-D) help recover more complete object templates.
@@ IV-A Pre-change 3DGS training
We first train a 3DGS model $\mathcal{G}$ on the pre-change images $\{I_{m}\}_{m=1}^{M}$. Following the standard 3DGS training procedure [8], we use a structure-from-motion (SfM) algorithm [33] to estimate pre-change camera parameters and generate a sparse point cloud, initializing 3D Gaussians for subsequent optimization (i.e. training). 111 For the success of SfM, we typically need the pre-change images to sufficiently cover the pre-change scene. But this requirement can be relaxed with sparse-view 3DGS training techniques, e.g. InstantSplat [34].
@@ IV-C EfficientSAM embedding for 2D change detection
We first coarsely detect 2D changes with EfficientSAM [11] image embeddings (Fig. 2). The SAM [23] image embedding, which encodes rich semantic information, has been shown to enable robust zero-shot change detection [9, 35]. However, probing SAM’s latent space can be computationally expensive. We therefore adopt the lightweight EfficientSAM encoder to extract image embeddings on the rendered and captured post-change images, and threshold their cosine similarity maps to obtain the change masks:
@@
As shown in Fig. 2, the high-confidence 2D segments on the rendered images are used as move-out masks (previous location) for the re-arranged objects, whereas the high-confidence masks on the captured images are identified as object move-in masks (new location):
@@ IV-D Change-aware object association
Starting from the object move-out masks on the first view, we iteratively associate the segments across remaining views by matching the partial point clouds. Specifically, we use the Hungarian algorithm to perform nearest-neighbor matching between the current object point cloud and the next segment with the Hausdorff distance. The matched new segments are added to the object templates while the unmatched segments are used to initialize new object templates. We also verify the semantic similarity of the matched segments by comparing their EfficientSAM embeddings. If their cosine similarity is below 40%, a new object template is created instead for the new segment. After the association is done, we obtain the object templates as the union of multi-view point clouds and features:
@@ IV-F Object pose change estimation
For moved objects, we use the 2D-3D correspondences from Sec.IV-D, and apply the standard RANSAC-PnP algorithm to coarsely estimate the object pose changes. The initial estimates are further optimized in Sec. IV-G.
@@
Figure 3: Qualitative results: Estimated 3D object masks projected on novel views. We show the projected object move-out (previous location) and move-in (new location) masks on the pre-change 3DGS render (left) and the post-change capture (right) from the same novel viewpoint for each scene. Please check out the supplementary video for mask projections on more views.
(a) Eval image
(b) Ground truth
(c) Ours
(d) MTP[19]
@@ V-C Baselines
D-NeRF-Diff [40]: Another baseline by C-NeRF authors that trains a D-NeRF with the two image sets as two time steps. It computes change masks using the absolute color differences between evaluation-view D-NeRF renders.
4DGS-Diff [29]: Similar to D-NeRF-Diff, we train 4DGS at two time steps and compute color differences between eval-view renders at different times to get the change masks.
On the 3DGS-CD dataset, we use MTP as baseline to predict change masks on rendered and captured images at evaluation views  [^3] C-NeRF is not used as a baseline on 3DGS-CD since it fails to train a post-change NeRF with reasonable quality using the sparse new images.. We adopt the ViT-B+RVSA model fine-tuned on the CDD dataset [41] as it performs best on our data. On the NeRF-CD dataset, we compare our method with C-NeRF, C-NeRF-Diff, D-NeRF-Diff and 4DGS-Diff.
@@ V-D Results
Our method is more accurate than the MTP baseline by a large margin, despite having a slower runtime due to the inference efficiency of MTP’s neural network. While MTP’s prediction accuracy could be improved through fine-tuning on in-domain data, our method exhibits strong zero-shot performance without any fine-tuning.
@@
Our method also outperforms the dynamic NeRF/3DGS-based baselines (i.e. D-NeRF-Diff, 4DGS-Diff) by a large margin in change detection accuracy. This is expected since the dynamic NeRF/3DGS models are not designed or optimized for predicting accurate 3D changes. We do not include a runtime performance comparison with these baselines because they don’t have an explicit change detection step. Comparing our method’s runtime to, for instance, the training time of 4DGS, would not be a fair comparison.
TABLE II: Quantitative evaluation of our method on the test scene Cats from the NeRF-CD [10] dataset. Runtimes for the *-Diff baselines are omitted as they don’t have an explicit change detection step.
@@ VI-A Object removal as prompt for object reconstruction
To illustrate this, we use the SOTA prompt-based 3DGS segmentation method SAGA [42] to isolate a target object from its 3DGS-represented scene, as shown in Fig. 6. While SAGA explicitly accounts for multi-level granularity, it still struggles to segment the entire object with a single click prompt ( or ), even with the best thresholds. We had to provide multiple precise clicks ( ) to extract the complete object. However, providing such precise clicks can be difficult on mobile devices. Similarly, using a language prompt generated by GPT-4o also failed to isolate the whole object.
To address this, we propose using object removal as an unambiguous prompt, as shown in Fig. 7. After capturing images of the object in its scene, users only have to remove the object, and take a few more images for the object-removed scene. Our method can then consistently identify the target object from the pre-change images, allowing us to train an object-3DGS without ambiguity. Compared to the click or language prompts, the object removal prompt is more intuitive and requires minimal user efforts.
@@ VI-B Robot workspace reset
Figure 8: 3DGS-CD enables robot workspace reset. Please click on the sub-figure (c) for the robot reset video.
@@
The robot workspace reset process can be summarized as follows: (1) RGB video scan to capture the initial state of the robot workspace; (2) Move objects in the workspace; (3) Capture a few more RGB images for the changed state of the workspace; (4) Use our method to estimate the moved objects’ shapes and pose changes in under 30s; (5) Robot pick and place to reset the state of the scene.
Key advantages of our robot workspace reset method over existing solutions (e.g. SayCan [43]) include: (1) No depth sensor or depth estimator is required. (2) No need for object class, model or object detectors. (3) No need for potentially ambiguous language prompts – Language prompts for a workspace with multiple identical objects present could confuse many existing models (e.g. SayCan [43]). (4) Precise reset of the object position and orientation.
Figure 8 shows a simulated experiment we conducted in PyBullet [44], where a robot arm resets the pose of a mug on a table. In this experiment, we used forward kinematics, rather than structure-from-motion (SfM) or visual localization, to estimate the pre- and post-change poses of the robot’s RGB camera. Without the SfM point cloud, the pre-change 3D Gaussians are randomly initialized before optimization. Once the middle mug’s pose change and shape are computed, we determine the pick-and-place poses for the robot gripper and use inverse kinematics to compute the joint motions. For simplicity, we use the initial gripper orientation as the grasp orientation and use the object’s centroid as the grasp position.
@@ VI-C Fast sparse-view-guided 3DGS update
In our method, we transform the in-object pre-optimized Gaussians according to the estimated object movements, and freeze all pre-trained Gaussian parameters. We then duplicate the Gaussians near the newly un-occluded regions and optimize only the duplicates to fill in the missing parts. Since these Gaussians are usually similar in position and color to the missing ones, they provide a good initialization for subsequent 3DGS optimization under sparse-view supervision. The Gaussians are optimized for 1000 step, during which we use the adaptive control strategy [8] to split, duplicate, reset and cull Gaussians every 100 steps.
We test our method on the Soda scene from the NeRF-Update dataset [9], as shown in Fig. 9. In this scene, 174 pre-change images are used to train a 3DGS and 4 post-change images serve as guidance for the update. We evaluate our method on reconstruction quality for the soda can’s move-in and move-out regions with PSNR and SSIM, and also on runtime performance, as reported in Tab. III. Our method achieves comparable reconstruction quality and superior runtime performance compared to the NeRF-update [9] method. We also compare it with the SOTA sparse-view 3DGS training method, InstantSplat [8], which is only trained on the 4 post-change images. While InstantSplat is faster, it delivers lower reconstruction quality.
@@ VII Limitations
Foreground-background semantic similarity: The success of our 2D change detection module depends on the semantic distinction between the moved object and its background. If the object and the background are semantically similar, such as removing an apple from a pile of apples, the change detection performance may degrade. This issue can be mitigated by integrating our EfficientSAM-based approach with a photometric differencing method, allowing for the assessment of both semantic and visual differences between the foreground and background.